Transit to Site 1127

The 23-nmi transit to Site 1127 was completed in 2.5 hr at 9.2 kt. A beacon was deployed at 1800 hr on 30 October, initiating operations at Site 1127. A second beacon was dropped to provide another working beacon for shallow-water operations, but it produced an erratic signal and was subsequently observed floating away, apparently self-released. A third beacon was deployed and the ship was stabilized on position.

Hole 1127A

An APC/XCB BHA was run to the seafloor, and Hole 1127A was spudded at 2110 hr on 30 October, recovering 9.68 m. Core 1H was curated, but the hole was terminated because recovery was inappropriate for establishment of a good mudline. The bit was repositioned for another mudline core, ending Hole 1127A at 2130 hr on 30 October.

Hole 1127B

The ship was not moved, and Hole 1127B was spudded at 2130 hr on 30 October. The bit was positioned at 487 mbrf, and Core 1H recovered 5.86 m, which indicated a water depth of 479.3 mbsl. APC coring advanced to 148.4 mbsf, orienting Cores 3H–16H. A noticeable odor of H2S was detected in Core 1H. H2S concentrations were first measured with a handheld device on Core 4H, which reached the maximum capability (2000 ppm) of the instrument. Core 6H (46.9 mbsf) had H2S concentrations of 94,953 ppm (gas chromatograph) measured from a gas void in the core (vacutainer). Coring operations were suspended for 2.5 hr while H2S safety procedures were implemented. These procedures included prepping crews, deploying safety equipment (e.g., Scott air packs), posting notices, installing curtains and a suction hose in the laboratory cutting room, deploying fans and fixed sensors, and increasing interior laboratory ventilation. The ship heading was rotated to maintain maximum air flow over the core handling area and air suction for the accommodation. Once a section was judged safe to enter the laboratories (Foss and Julson, 1993), it was run through the multisensor track (MST), split, and allowed to degas under the vacuum in the splitting room. Ten more cores were taken with little entry of cores into the laboratories. Coring was suspended again to determine an improved method for more rapid degassing of the cores. Sections were run through the MST, split, carried out to the catwalk, and opened there to allow safer and more rapid venting. When the super saw was required to cut cores, sections were split and dressed out in the splitting room and then carried outside for continued venting. APC coring continued thereafter on a core-by-core basis. The coring rate had to be slowed to maintain proper handling and racking of the core, recharging of air packs, and cleaning and airing of core to reduce the ambient H2S level.

Core 16H had 60K lb overpull with high H2S concentrations. Therefore, coring was switched to the XCB system for the potential reduction in gas volumes typically observed in switching to XCB coring. Cores 17X and 18X were cut from 148.4 to 164.5 mbsf with good recovery; however, H2S readings remained ~105 ppm. Coring continued with the XCB from 164.5 to 510.7 mbsf (Cores 19X–55X) on a core-by-core basis. Experience proved that coring more rapidly than ~45 min/core exceeded our capacity to process the core on the catwalk while wearing full protective gear.

The hole was prepared for logging with a wiper trip to 100 mbsf, and the bottom was tagged at 510.7 mbsf with no fill. The hole was filled with 180 bbl of sepiolite mud to protect logging tools from potential H2S exposure, and the logging cable was coated with a 50/50 mixture of Marvel Mystery Oil and diesel. The bit was pulled up to the logging depth at 87 mbsf. The triple combo (without radioactive sources), FMS-Sonic, and GHMT logs were run to 510.7 mbsf. The WST was not required because in this thick, uniform sequence, the Sonic log proved to be adequate. The hole was plugged with 40 bbl of 10.5-ppg mud, and the pipe was pulled. The rig was secured for transit at 0230 hr on 3 November, ending Site 1127.

To 182 Summary of Engineering and Drilling Operations for Site 1128

To 182 Table of Contents